The probabilities of a boy passing English and Mathematics tests are x and y respectively. Find the probability of the boy failing both tests.

1-(x-y)+xy
1-(x+y)-xy
1-(x+y)+xy
1-(x-Y)-xy
1
Expert's answer
2015-02-25T13:10:48-0500
The probability of the boy failing both tests is (1-x)(1-y)=1-y-x+xy=1-(x+y)+xy.